OpenStudy (anonymous):

Two boxes of apples contain a total of 24 apples. If you halve the number of apples in the first box and add 4 apples to the second box, the total changes to 20 apples. How many apples are in each box initially?

OpenStudy (anonymous):

are you supposed to use "systems of equations"? or can you do anything?

OpenStudy (anonymous):

I'm not sure... This is an online pre test I'm taking and it's just kind of like "Here you go, do this." ;_;

OpenStudy (anonymous):

if it was me, since the answer is obviously two whole numbers, i would guess and check for example lets start with 12 in each take half of 12, get 6 add 4 to 12, get 16 but 6+16=22, so that is not right how about 14 and 10?

OpenStudy (anonymous):

that doesn't work either, because half of 14 is 7 and 10 + 4 = 14 and \(7+14=21\) i bet we get it in the next try 16 and 8

OpenStudy (anonymous):

half of 16 is 8, 8 +4 = 12 and 8+12=20 done

OpenStudy (anonymous):

Ok, thank you so very much. ;3; I'm really no good in mathematics, and not even my teacher can help me. ;3;